VEGAS Team Meeting

Dial in Number: 877-937-8961

Passcode: 9445091#

Attendees: John, Randy, Adam, Joe, Sam, Mark, Paul, Marty, Dave, Jason

Agenda:

Whole group:

  1. Status of BOFs for Modes 2 and 3 (H16k)

GB only

  1. Tsys stability
  2. FTarget values
  3. Subbeam Nod
  4. Deltafreq
  5. Modes 14/19
  6. fcenter shift
  7. non-constant baselines
  8. Plans for test time

Meeting notes:

Whole group:

  1. Status of BOFs for Modes 2 and 3 (H16k)
    • Waiting BOF rebuild from Hong, and testing/calibration from Joe.
    • John to make sure Xylinx license is current after SOC server outage.

GB only

  1. Tsys stability
    • Tsys is higher during the first integration (GBTIDL high by afactor of two)
      • Workaround is to eliminate first integration
      • Sam & Richard to look at fits issues while Joe looks at deeper fits data
      • Research if the counts are what is off by factor of two
  2. IFTarget values
    • Need to verify whether the IF Target values set automatically are adequate for both wide bandpasses and narrow bandpasses when used with VEGAS.
    • hypotheses: must be levels between IF rack --> converter rack -->VEGAS
    • Function of signal density on fiber - need to adjust to stay in range of ADCs
    • Since VEGAS is wideband, signals should be set to the levels correct for wideband observations
    • This is a GBT system problem - not VEGAS. Adam will do some converter rack tests and assemble results documentation for presentation to observing support scientists
  3. Subbeam Nod
    • Manager possibly not configured for nod or subbeam nod?
    • This has not yet been tested w/VEGAS and potentially not any different than the Spectrometer
    • Conduct tests w/VEGAS
    • If tests show any issues, shorter integrations may be mitigation
  4. Deltafreq
    • Not fully implemented or tested
    • Another system problem, not seen as a VEGAS problem (and possibly a Spectrometer problem too) - non-VEGAS scientists should investigate
  5. Modes 14 & 19
    • High resolution modes not yet available
    • Randy is working issues and looks hopeful with fix
    • Medium priority hardware system problem
  6. fcenter shift
    • Center frequency still shifted to avoid the center channel spike,to remove this?
    • This will be with us always (interleave/clock frequency artifact)
    • This needs to be documented and added to Observing instructions
    • This is handled by interpolation code in SDFITS (Final solution)
  7. non-constant baselines:
    • VEGAS baselines are not constant, and fluctuate a bit in offon observing (see here - The red is from the spectrometer, the black from VEGAS). Looking at two 'Offs' by way of (off1-off2)/off2 - the way they would be calibrated as OffOn scans (see below). These are not perfect, but not too terrible, possibly suggesting a non-linearity => iftarget issues?
    • off-on:
      off off.png
    • Retest "On-On" (mode 4) at Q band to see output without so much RFI
  8. Plans for test time
    • Unscheduled request(s):
      • new BOFs when available (Prestage/Sam/Brandt)
      • IF rack & Converter rack level tests (Kobelski)
      • Subbeam nodding tests (Frayer & Brandt)
      • Retest "On-On" (mode 4) with noise source and then at Q band to see output without so much RFI (Prestage?)
    • Tuesday
      • pm IF rack & Converter rack level tests (Kobelski)
    • Wednesday
      • PM IF rack & Converter rack level tests (Kobelski)
    • Thursday
      • Subbeam nodding tests (Frayer & Brandt)
      • Deep integration tests w/mode 2 and new BOFs (Frayer)
    • Monday (8/26)
      • None at this time

Other

  1. new meeting time - 11:00 EDT Mondays
  2. Create regression test scripts for M&C plus VEGAS
  3. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-08-18
Topic revision: r2 - 2014-08-26, AdamKobelski
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback